# Cell bulging and extrusion in a three-dimensional bubbly vertex model for curved epithelial sheets

### v1

@@ II.3 Shift of icosahedral buckling threshold
Figure 2: Lowered buckling threshold for icosahdral instability in the bubbly vertex model (BVM) compared to the vertex model (VM). (a,b) Icosahedral shells with Caspar Klug indices $(8,0)$ with tension $\Gamma=0.6$ in the VM (a) and the BVM (b). (c) Asphericity $\alpha=\langle(R-\langle R\rangle)^{2}\rangle/\langle R^{2}\rangle$ of the cell centers in the BVM shell (symbols) and of the continuum curve describing the VM (taken from Refs. [19, 18]). Inset shows shell with size $(2,0)$, $\Gamma=0.6$. Radii are rescaled by the Young’s modulus $Y$, bending rigidity $\kappa$ and a nonlinearity correction $k_{\mathrm{ico}}$, see Appendix B. (d) For a small shell (Caspar Klug indices $(2,0)$) the average ratio of luminal and outer interfacial area in the BVM (solid) and VM (dashed). Inset shows pentagonal cell with collapsed luminal interface ($\Gamma=1.0$). Pentagonal cells are shown in red.
@@
$\alpha=\langle(R-\langle R\rangle)^{2}\rangle/\langle R^{2}\rangle,$ (2)
[…]
We also find that the asphericities do not fall on one curve, but rather deviate from an average curve, which results from the broad distributions of hexagonal opening angles described above and therefore less regular structures. Similarly to the VM, we have a collapse of the luminal interface for larger $\Gamma$, which is much more pronounced in the BVM with a stronger deviation of the luminal-to-outer area ratios between pentagonal and hexagonal cells for small spheres, cf. Fig. 2 (d). Even shells as small as $(2,0)$, i.e. only one hexagonal cell between the pentagonal cells, show a big difference between pentagonal and hexagonal cells, and thus adopt a strongly icosahedral shape. We thus see that the buckling threshold is reduced to effectively one neighboring cell in the BVM, rendering the icosahedral tissue-scale instability much more relevant in the case of curved cellular interfaces in the BVM.
III Continuum theory of a defect cell in a mean-field tissue
@@ III.1 Parameterizing the defect cell
To understand the nature of the significant lowering of the icosahedral buckling threshold for the BVM compared to the VM, we now investigate the BVM in a continuum setting. We focus on the icosahedral tip and the conical instability to first develop the continuum model in a setting amenable to comparison to numerical simulations. At the site of the disclination defect we excise the continuous material and directly consider the energy of the pentagonal cell, following the description of the bubbly vertex model, cf. Eq. (1). Fig. 3(a) depicts the geometry we assume.
Figure 3: Schematic depiction of the defect cell in a mean field tissue. (a) The tissue profile is shown as a radial function in cylindrical coordinates. At the center is the defect cell, which consists of a base of a cone with spherical caps on the apical and basal sides. This defect is surrounded by mean-field tissue that forms a spherical cap segment (light grey) and for larger distance a cone (dark grey). (b) The model can describe different defect cell geometries, namely, the vertex-model reference case ($h_{S}^{a}=h_{S}^{b}=0$), a completely bulged case ($r_{b}=0$, $\delta/2>0$) and a fully extruded case ($r_{b}=0$, $\delta/2=0$). (c) For pentagonal cells we consider a geometrical correction, where we assume a pentagonal pyramid base with two spherical caps. The radii are chosen such that the area of the caps’ disc bases match the pentagonal areas (center). and the slant height $\tilde{h}$ is considered at the lateral faces’ centers.
@@
With the chosen parameterization we are able to describe different defect cell morphologies, i.e. a vertex model case with flat apical and basal interfaces, a bulged case with a collapsed interface, and an extruded case (Fig. 3(b)). Due to the pentagonal nature of the defect, we modify the conical shape to reflect the pentagonal structure by considering a pentagonal pyramid base, see Fig. 3(c). For this we introduce a geometrical factor $\zeta_{\mathrm{geom}}$, which relates the basal and apical polygonal edge lengths, $b$ and $a$, respectively, to the areas such that for the apical area $A_{\mathrm{a}}=\zeta_{\mathrm{geom}}a^{2}$ and similarly for the basal side. Additionally, we introduce $\eta_{\mathrm{geom}}$, which relates the radius of a regular polygon’s incircle $r_{a,b}$, i.e. the largest circle that fits into the regular polygon, to the edge length, e.g., for the apical side $r_{a}=\eta_{\mathrm{geom}}a$.
For a bulged pentagonal cell we consider the pyramid base with apical/basal spherical caps. We assume the caps to have the same disc area as the apical/basal pentagons, neglecting the slight mismatch of the two different shapes, cf. Fig. 3(c). This is consistent with simulated cell shapes in the BVM as shown in Fig. 1(a). For the spherical caps we denote the base disc radii with $t_{(\cdot)}$. The cap’s base area is $\pi t_{(\cdot)}^{2}$. This implies $\pi t_{a}^{2}=\zeta_{\mathrm{geom}}a^{2}$ for the apical side and analogously for the basal side. The radius of the apical spherical cap’s base then relates to the edge length via $t_{a}=\sqrt{(\zeta_{\mathrm{geom}}/\pi)}\,a$ and analogously for the basal side.
Henceforth we denote the side to which the tissue buckles (positive $z$) the basal side and the opposite side the apical one, as this corresponds to the sign of curvature at the villus tip in the small intestine. The opening angle of the pyramid at the lateral face centers, $\delta/2$, the basal inradius $r_{b}$ and the tissue height $\tilde{h}$ define the base part. To consider the energy change due to extrusion and bulging, we define as the reference case the vertex model case, where the defect cell is not bulged and induces the same curvature as the tissue would. The spherical caps of the defect cell are uniquely defined by the basal and apical radii and the height of the caps ${h_{S}^{a,b}}$ (cf. Fig. 3).
The radius of curvature of the spherical caps $R_{S}^{a,b}$ depends on the pressure differences between the cell and the surrounding medium $\Delta P_{a,b}$ at the corresponding side via the Young-Laplace law
$\Delta P_{a,b}=\frac{2\Gamma_{a,b}}{R_{S}^{a,b}}.$ (3)
@@
For the sake of simplicity we assume $\Gamma_{\mathrm{a}}=\Gamma_{\mathrm{b}}$ in this work, i.e., no apico-basal polarity. Using this relation, we express the basal cap height via the apical one, assuming that extrusion only occurs on the apical side and therefore choosing the smaller of two possible cap heights.
@@
The volume of the cell then implicitly defines $h_{S}^{a}$ via
$1=\frac{\tilde{h}\cos(\delta/2)}{3}\zeta_{\mathrm{geom}}\left(a^{2}+ab+b^{2}\right)\\ +\frac{h_{S}^{b}}{6}\left(3\zeta_{\mathrm{geom}}b^{2}+\pi{h_{S}^{b}}^{2}\right)+\frac{h_{S}^{a}}{6}\left(3\zeta_{\mathrm{geom}}a^{2}+\pi{h_{S}^{a}}^{2}\right),$ (5)
@@
with the lateral (slant fraction) height $\tilde{h}$. The last two terms correspond to the spherical caps and vanish for the non-bulged reference cell as $h_{S}^{b}=h_{S}^{a}=0$. The apical and basal sides are related to each other via the opening angle, $a=b+\sin(\delta/2)\tilde{h}/\eta_{\mathrm{geom}}$, cf. Fig. 3(c). In the deformed cell, which we will now denote with a prime, the cap heights are then determined by $b^{\prime},\delta^{\prime}$ and $\tilde{h}$.
@@
In a continuum approach we would not resolve the defect cell and the surrounding spherical cap with spherical radius $R_{0}$ would determine the opening angle, i.e., $\sin(\delta/2)=\eta_{\mathrm{geom}}\left(1+\frac{R_{0}+\tilde{h}/2}{R_{0}-\tilde{h}/2}\right)b/2R_{0}$. However, resolving the defect cell explicitly allows us to consider a vertex model later, which goes beyond this approximation, even if we do not allow for interfacial curvature.
@@
With known cap heights we can write down the energy of the bubbly defect cell by adding the lateral surface energies and the surface energies of the spherical caps to obtain, where the cell is given by a regular $n$-gon (i.e. we first consider $n=5$),
$E_{\mathrm{core}}^{\mathrm{bulged}}=\Gamma_{a}\left(\zeta_{\mathrm{geom}}a^{2}+\pi{{h_{S}^{a}}}^{2}\right)+\Gamma_{b}\left(\zeta_{\mathrm{geom}}b^{2}+\pi{{h_{S}^{b}}}^{2}\right)\\ +\frac{n\tilde{h}}{4}\left(2b+\frac{\tilde{h}}{\eta_{\mathrm{geom}}}\sin(\delta/2)\right).$ (6)
@@
For a pentagonal defect we have $\zeta_{\mathrm{geom}}=\sqrt{5(5+2\sqrt{5})}/4$ and $\eta_{\mathrm{geom}}=1/\left(2\sqrt{5-\sqrt{20}}\right)$. We assume that the lateral slant height of the pyramid base is the equilibrium height of the surrounding tissue $\tilde{h}$ and thus the only two free parameters are $b$, $\delta$. To describe the full tissue we additionally need the background curvature $R_{0}$, which relates to the buckling radius $R_{B}$ via the conical opening angle $\gamma$, cf. Fig. 3(a).
[…]
@@
Fig. 4 shows the dimensionless core energy, Eq. (6), for different opening angles $\delta/2$ as a function of the basal edge length $b$ for (a) $\Gamma=\Gamma_{\mathrm{a}}=\Gamma_{\mathrm{b}}=0.65$ and (b) $1.0$, respectively. Edge length $b$ is bounded from above by the volume constraint in the unbulged vertex model case with $h_{S}^{b}=h_{S}^{a}=0$. For each angle the energy initially decreases if we decrease $b$ with concomitant outward bulging (top panel), as the area to volume ratio decreases. For even smaller $b$ the energy usually increases again, suggesting that a configuration with curved interfaces is always energetically more favorable.
@@
Increasing the angle leads to an increase in the core energy in the vertex model case (cf. dashed line at which the curves end in the top panel) and shifts the minimum of the energy toward smaller $b$ if we allow for bulging. If we compare the vertex model energy with the minimal energy in the bubbly vertex model for given opening angles (and thus surrounding tissue curvature), we find a smaller energy in the latter case, see lower panel.
@@ III.2 Deformation of the mean-field tissue
We use the description as a moderately bent plate as derived in Ref. [18]. The energy change due to the defect deformation can then be split into different contributions. First, the bending energy will change, both for the mean and Gaussian curvature contributions, due to changes in the curvature radii and opening angles in the inner cap segment. Second, the stretching energy will change as the azimuthal strain stemming from the pentagonal defect can be relaxed through changes in the curvature. For simplicity we will neglect radial stretching contributions at the interfaces by considering stress free boundary conditions at the contact points of defect cell, cap segment and outer cone segment and at the outer edge of the cone, which we will justify numerically in the following.
@@
III.2.1 Bending energies
For the cap segment we consider the radial deflection function $f$ with rotational symmetry
$f(r)=R_{0}-\sqrt{R_{0}^{2}-(\rho+r)^{2}},$ (8)
$e_{\mathrm{bending}}=\frac{\kappa}{2}H^{2}+\kappa_{\mathrm{G}}K,$ (9)
$H(r)$ $=\frac{1}{r}\partial_{r}r\left(\frac{\partial_{r}f}{\sqrt{1+(\partial_{r}f)^{2}}}\right)=\frac{2}{R_{0}}+\frac{\rho}{rR_{0}},$ (10)
$K(r)$ $=\frac{\frac{1}{r}(\partial_{r}f)(\partial_{r}^{2}f)}{\left(1+(\partial_{r}f)^{2}\right)^{2}}=\frac{\rho+r}{rR_{0}^{2}}.$
$\frac{E_{\mathrm{bending}}^{\mathrm{mean}}}{\kappa\pi}=-\frac{4}{R_{0}}\sqrt{R_{0}^{2}-(\rho+r)^{2}}-\frac{\rho^{2}}{\sqrt{R_{0}^{2}-\rho^{2}}}\times\\ \left.\mathrm{artanh}\left(\frac{R_{0}^{2}-\rho\left(\rho+r\right)}{\left(R_{0}^{2}-\rho^{2}\right)^{1/2}\left(R_{0}^{2}-\left(\rho+r\right)^{2}\right)^{1/2}}\right)\right|_{r=r_{0}}^{r=R_{B}}.$ (11)
Note that for $\rho=r_{0}=0$ and small buckling radius $R_{B}$, i.e. small $|\nabla f|$ in the outer cone segment, the Taylor approximation in $R_{B}$ scales quadratically, consistent with the approximation of the total energy with a flat disc as proposed by Seung and Nelson [22] for the flattened tip for the deformed shape around the defect. The first term in Eq. (11) arises from the spherical contribution of the mean curvature bending energy, cf. corresponding first term in the mean curvature in Eq. (10). The second term, which is proportional to the squared radial offset $\rho^{2}$, serves as a correction for reducing the radius of the circular segment in cylindrical coordinates, as we decrease the radius of curvature in the azimuthal direction via this offset. This term stems from the quadratic contribution of $\rho^{2}$ from the squared mean curvature, Eq. (10), in the bending energy.
with tissue opening angle $\gamma/2$, cf. Fig. 3(a).
$E_{\mathrm{cone}}=\frac{f_{1}}{(1+f_{1}^{2})^{1/2}}\pi\kappa\log(R/R_{B}),$ (13)
with the total radius $R$, the buckling radius $R_{B}$, and the cone slope
$f_{1}^{\mathrm{cone}}=\left[\frac{1}{(1-s/2\pi)^{2}}-1\right]^{1/2}.$ (14)
III.2.2 Stretching energies
The stretching energy density reads
$\varepsilon_{rr}$ $=w^{\prime}(r)-1+\frac{(f^{\prime}(r))^{2}}{2},$ (16)
$\varepsilon_{\theta\theta}(r)$ $=\frac{w(r)-r}{r}+q_{s},$
$\varepsilon_{r\theta}$ $=0,$
We now want to take a different approach, however, and consider a ring of material with inner and outer radii $r_{0}$ and $R_{B}$, deflected with a known fixed slope $f^{\prime}=f_{1}$. Introducing $\Delta s=q_{s}-{f_{1}}^{2}/2$, Eq. (17) is solved by
$w(r)=\frac{\mu}{2\mu+\lambda}\Delta sr\log(r)+\frac{k_{1}}{r}+k_{2}r,$ (18)
$\Delta s=q_{s}-\left(1-\frac{1}{\sqrt{1+{f_{1}}^{2}}}\right).$ (19)
As boundary conditions we assume vanishing radial stresses at the inner and outer radii, i.e. at $r=r_{0}$ and $r=R_{B}$
This boundary condition assumes that no radial (and thus total compressive) stresses act on the boundary of the cap segment and the conical part of the surrounding tissue. Comparing the cell height in VM simulations to the theoretical mean-field tissue height serves as an estimate for the compressive in-plane stress. The height difference from the mean-field height (cf. Appendix B) is shown in Fig. 5(a).
Figure 5: Comparison of continuum model with vertex model simulations of icosahedral shells. (a) VM shell with Caspar-Klug size $(20,0)$ and $\Gamma=1.1$ with top view and side view cross-section. Color-coded is cell height difference from the mean-field VM height (cf. Eq. (32)). (b) Total energy of the proposed full defect model as function of defect opening angle $\delta/2$ for different apical/basal surface tensions $\Gamma$. The dotted grey line indicates the angle boundary from the volume constraint. (c) Opening angles $\delta/2$ (top) and basal to apical face area ratios (bottom) for the different models. The defect model (solid) considers a defect in a mean-field tissue with the angle $\delta/2$ corresponding to the minimal energy configuration, cf. (a). The spherical cap model (dashed) corresponds to this energy under the assumption that $\delta/2$ matches the background curvature. The symbols indicate vertex model simulation results for different Caspar-Klug indices $(T_{1},T_{2})$.
$\frac{E_{\mathrm{stretch}}}{\pi Y\left(1+f_{1}^{2}\right)^{1/2}}$ $=\frac{\Delta s^{2}}{8}\left(R_{B}^{2}-r_{0}^{2}\right)$ (21)
$\ -\frac{\Delta s^{2}}{2}\ \frac{R_{B}^{2}r_{0}^{2}}{R_{B}^{2}-r_{0}^{2}}\log^{2}\left(\frac{R_{B}}{r_{0}}\right).$
Finally, we can write down the full energy of the surrounding tissue as a sum of the contributions in Eqs. (11, 12, 13, 21), i.e.
III.3 Comparison to icosahedral vertex model shells
The corresponding continuum model energy for this vertex model case consists of the core energy for a defect cell without bulging apical/basal membranes, i.e., with $h_{S}^{a}=0=h_{S}^{b}$, Eq. (6); the bending energy of both the cap segment, Eqs. (11, 12), and the outer cone region, Eq. (13); and the stretching energy in the cap, Eq. (21). For the resulting total energy $E_{\mathrm{tot}}$ we assume a large outer radius of the total disc $R$, such that area changes from deformations in the cone are negligible, and thus drop the term proportional to $\log(R)$ in Eq. (13), as we look at energy differences with respect to a reference state.
In the derivation of the mean field tissue (cf. Appendix B) lateral interface energies in the energy density are included by counting them half for the respective cell area (to avoid double counting). Here we now introduce a boundary energy term, which subtracts this lateral energy from the mean field tissue and adds half the pentagonal lattice energy of the defect. This lateral energy is present in full and should be considered completely, due to the explicit description of the defect cell. For the lateral contact between the defect and the tissue we thus subtract the lateral energy contribution which is implicitly given in the mean-field tissue boundary (with a circular cut-out) and replace it by the actual lateral energy contribution of the cell (with a pentagonal shape). This energy reads
$E_{\mathrm{bndry}}=\frac{5\tilde{h}}{4}\left(2b+\frac{\tilde{h}}{\eta_{\mathrm{geom}}}\sin(\delta/2)\right)-\pi r_{0}\tilde{h}$ (23)
and it is added to the total energy, i.e.
$E_{\mathrm{tot}}=E_{\mathrm{core}}+E^{\mathrm{mean}}_{\mathrm{bending}}+E^{\mathrm{Gauss}}_{\mathrm{bending}}+E_{\mathrm{cone}}\\ +E_{\mathrm{stretch}}+E_{\mathrm{bndry}}.$ (24)
For the elastic parameters and the bending rigidities of the mean-field tissues we use the mean-field values derived for the VM in Ref. [18], i.e. Eqs. (33, 35) in Appendix B.
We found that for every $\delta/2$ and b the energy has a unique minimum as a function of the buckling radius $R_{B}$, or, equivalently, as a function of the radial cap curvature radius $R_{0}$. We thus minimize the energy with respect to $R_{B}$ to obtain the total energy as a function of the opening angle $\delta/2$ in the vertex model. Fig. 5(b) depicts the energies for different surface tensions $\Gamma$, where we observe a clear minimum. Note that for large $\Gamma$ the volume constraint yields an upper bound for the possible opening angles. Our approach, the defect model, allows the defect to have an opening angle independently from the background tissue curvature. For comparison we also consider the case of the opening angle following the background curvature, i.e., $\delta/2$ is chosen such that it matches $R_{0}$ with $\rho=0$, which we henceforth denote as the spherical cap model. The offset of the energies is a consequence of the neglection of the $\log(R)$-term and results from the different cell number densities for different $\Gamma$.
IV Mechanisms of cell extrusion
@@
IV.1 Extrusion of a bubbly defect cell in a mean-field vertex model tissue

### v2

@@ II.5 Shift of icosahedral buckling threshold
Figure 5: Lowered buckling threshold for icosahedral instability in the bubbly vertex model (BVM) compared to the vertex model (VM). (a,b) Icosahedral shells with Caspar Klug indices $(8,0)$ with tension $\Gamma=0.6$ in the VM (a) and the BVM (b). (c) Asphericity $\alpha$ in the BVM-shell (symbols) and of the continuum curve describing the VM (taken from Refs. [20, 19]). Inset shows shell with size $(2,0)$, $\Gamma=0.6$. Radii are rescaled by the Föppl-von Kármán number and with a correction $k_{\mathrm{ico}}$ from non-linear elasticity. (d) For a small shell (Caspar Klug indices $(2,0)$) the average ratio of luminal and outer interfacial area in the BVM (solid) and VM (dashed). Inset shows pentagonal cell with collapsed luminal interface ($\Gamma=1.0$). Pentagonal cells are shown in red.
@@
$\alpha=\frac{\langle(R-\langle R\rangle)^{2}\rangle}{\langle R^{2}\rangle},$ (4)
where $R$ is the radial distance of a cell (determined as the average of the centers of mass of the apical and basal interfaces) from the shell center and $\langle\cdot\rangle$ is the multicellular average. For the VM, this quantity has been shown before to follow a continuous curve as a function of the rescaled shell size, corresponding to the Föppl-von Kármán number $\gamma_{\mathrm{FvK}}\equiv Y_{\mathrm{eff}}R^{2}/\kappa$, with bending rigidity $\kappa$, radius $R$, and (effective) Young’s modulus $Y_{\mathrm{eff}}$, corrected here for effects from nonlinear elasticity with a numerical factor $k_{\mathrm{ico}}$ [19]. As shown in Fig. 5(c), the asphericity is much larger for small shells in the BVM compared to the VM. Details on the axis scaling are given in Appendix H.
We also find that the asphericities in the BVM do not fall on one curve, but rather deviate from an average curve, which results from both the broad distributions of hexagonal opening angles described above and therefore less regular structures, and from cellular positioning for small cell numbers on the icosahedral surface due to undersampling of the icosahedral surface. Similarly to the VM, we have a collapse of the luminal interface for larger $\Gamma$, which is much more pronounced in the BVM with a stronger deviation of the luminal-to-outer area ratios between pentagonal and hexagonal cells for small spheres, cf. Fig. 5 (d). Even shells as small as $(2,0)$, i.e. only one hexagonal cell between the pentagonal cells, show a big difference between pentagonal and hexagonal cells, and thus adopt a strongly icosahedral shape. We thus see that the buckling threshold is reduced to effectively one neighboring cell in the BVM, rendering the icosahedral tissue-scale instability much more relevant in the case of curved cellular interfaces in the BVM.
III Continuum theory for extrusion of a defect cell in a mean-field tissue
@@ III.1 Parameterizing the defect cell
To understand the nature of the significant lowering of the icosahedral buckling threshold for the BVM compared to the VM and expand the model to capture cell extrusion, we now investigate the BVM analtically in a continuum setting. We focus on the icosahedral tip and the conical instability to first develop the continuum model in a setting amenable to comparison to numerical simulations. At the site of the disclination defect we excise the continuous material and directly consider the energy of the pentagonal cell, following the description of the BVM, cf. Eq. (1). Fig. 6(a) depicts the geometry we assume.
Figure 6: Schematic depiction of the defect cell in a mean field tissue. (a) The tissue profile is shown as a radial function in cylindrical coordinates. At the center is the defect cell, which consists of a base of a cone with spherical caps on the apical and basal sides. This defect is surrounded by mean-field tissue that forms a spherical cap segment (light grey) and for larger distance a cone (dark grey). (b) The model can describe different defect cell geometries, namely, the vertex-model reference case ($h_{\mathrm{S}}^{\mathrm{a}}=h_{\mathrm{S}}^{\mathrm{b}}=0$), a completely bulged case ($r_{\mathrm{a}}=0$, $\delta/2>0$) and a fully extruded case ($r_{\mathrm{a}}=0$, $\delta/2=0$). (c) For pentagonal cells we consider a geometrical correction, where we assume a pentagonal pyramid base with two spherical caps. The radii of the caps, $\bar{r}_{\mathrm{a,b}}$, are chosen such that the area of the caps’ disc bases, the circles, match the pentagonal areas (center). and the slant height $\tilde{h}$ is considered at the lateral faces’ centers.
@@
With the chosen parameterization we are able to describe different defect cell morphologies, i.e. a vertex model case with flat apical and basal interfaces, a bulged case with a collapsed interface, and an extruded case (Fig. 6(b)). Due to the pentagonal nature of the defect, we modify the conical shape to reflect the pentagonal structure by considering a pentagonal pyramid base, see Fig. 6(c). For this we introduce a geometrical factor $\zeta_{\mathrm{geom}}$, which relates the basal and apical polygonal edge lengths, $b$ and $a$, respectively, to the areas such that for the apical area $A_{\mathrm{a}}=\zeta_{\mathrm{geom}}a^{2}$ and similarly for the basal side. Additionally, we introduce $\eta_{\mathrm{geom}}$, which relates the radius of a regular polygon’s incircle $r_{\mathrm{a,b}}$, i.e. the largest circle that fits into the regular polygon, to the edge length, e.g., for the apical side $r_{\mathrm{a}}=\eta_{\mathrm{geom}}a$.
For a bulged pentagonal cell we consider the pyramid base with apical/basal spherical caps. We assume the caps to have the same disc area as the apical/basal pentagons, neglecting the slight mismatch of the two different shapes, cf. Fig. 6(c). This is consistent with simulated cell shapes in the BVM as shown in Fig. 1(a). For the spherical caps we denote the base disc radii with $t_{\mathrm{a}}$; the cap’s base area is $\pi t_{\mathrm{a}}^{2}$; this implies $\pi t_{\mathrm{a}}^{2}=\zeta_{\mathrm{geom}}a^{2}$ for the apical side and analogously for the basal side. The radius of the apical spherical cap’s base then relates to the edge length via $t_{\mathrm{a}}=\sqrt{(\zeta_{\mathrm{geom}}/\pi)}\,a$ (analogously for basal).
Henceforth we denote the side to which the tissue buckles (positive $z$) the apical side and the opposite side the basal one, as this corresponds to the situation in the experimental mCol-organoids when considering closed shells. Note that the sign of curvature at the villus tip in the small intestine is opposite. The opening angle of the pyramid at the lateral face centers, $\delta/2$, the apical inradius $r_{\mathrm{a}}$ and the tissue height $\tilde{h}$ define the base part. To consider the energy change due to extrusion and bulging, we define as the reference case the vertex model case, where the defect cell is not bulged and induces the same curvature as the tissue would. The spherical caps of the defect cell are uniquely defined by the basal and apical radii and the height of the caps $h_{\mathrm{S}}^{\mathrm{a,b}}$} (cf. Fig. 6).
The radius of curvature of the spherical caps $R_{\mathrm{S}}^{\mathrm{a,b}}$ depends on the pressure differences between the cell and the surrounding medium $\Delta P_{\mathrm{a,b}}$ at the corresponding side via the Young-Laplace law
$\Delta P_{\mathrm{a,b}}=\frac{2\Gamma_{a,b}}{R_{\mathrm{S}}^{\mathrm{a,b}}}.$ (5)
@@
As described earlier, we assume $\Gamma_{\mathrm{a}}=\Gamma_{\mathrm{b}}$ in the analytical treatment, i.e., no apico-basal polarity. Using this relation, we express the basal cap height via the apical one, assuming that extrusion only occurs on the outer (basal) side (similar to the sign of curvature at the intestinal villus tip) and therefore choosing the smaller of two possible cap heights.
@@
The volume of the cell then implicitly defines $h_{S}^{a}$ via {multline} 1 = ~hcos(δ/2)3 ζ_geom (a^2 + ab + b^2)
@@
+ hSb6 ( 3 ζ_geomb^2 + πh_S^b^2) + hSa6 ( 3 ζ_geoma^2 + πh_S^a^2), with the lateral (slant fraction) height $\tilde{h}$. The last two terms correspond to the spherical caps and vanish for the non-bulged reference cell as $h_{\mathrm{S}}^{\mathrm{b}}=h_{\mathrm{S}}^{\mathrm{a}}=0$. The apical and basal sides are related to each other via the opening angle, $b=a+\sin(\delta/2)\tilde{h}/\eta_{\mathrm{geom}}$, cf. Fig. 6(c).
@@
In a continuum approach we would not resolve the defect cell and the surrounding spherical cap with spherical radius $R_{0}$ would determine the opening angle, i.e., $\sin(\delta/2)=\eta_{\mathrm{geom}}\left(1+\frac{R_{0}+\tilde{h}/2}{R_{0}-\tilde{h}/2}\right)a/2R_{0}$. However, resolving the defect cell explicitly allows us to consider a vertex model later, which goes beyond this approximation, even if we do not allow for interfacial curvature.
@@
With known cap heights we can write down the energy of the bubbly defect cell by adding the lateral surface energies and the surface energies of the spherical caps to obtain, where the cell is given by a regular $n$-gon, {multline} E_core^bulged = Γ_a ( ζ_geom a^2 + πh_S^a^2) + Γ_b ( ζ_geom b^2 + πh_S^b^2)
@@
+ n~h4 (2a + ~hηgeom sin(δ/2)). For a pentagonal defect we have $n=5$, $\zeta_{\mathrm{geom}}=\sqrt{5(5+2\sqrt{5})}/4$, and $\eta_{\mathrm{geom}}=1/\left(2\sqrt{5-\sqrt{20}}\right)$. We assume that the lateral slant height of the pyramid base is the equilibrium height of the surrounding tissue $\tilde{h}$ and thus the only two free parameters are $a$ and $\delta$. To describe the full tissue we additionally need the background curvature $R_{0}$, which relates to the buckling radius $R_{\mathrm{B}}$ via the conical opening angle $\gamma$, cf. Fig. 6(a).
Figure 7: The core energy for the different defect cell configurations for (a) small apical/basal surface tension $\Gamma$ and (b) large $\Gamma$. The solid colored curves in the upper panels are the energies as a function of apical edge length $a$ for fixed opening angles $\delta/2$. The curves end at the vertex line (dashed) at which $a$ is maximal for the angle as such that the apical and basal faces do not bulge. All possible states and energies which can be reached with $\delta/2$ and $a$ such that the volume constraint is fulfilled are shaded in grey. The minimum line (dotted) indicates the minimal energy for given $a$. In the lower panels the vertex line and the minimum line are parameterized with $\delta/2$. The minimum line is obtained by finding the edge length $a_{\mathrm{min}}$ that minimizes $E_{\mathrm{core}}$ for a given $\delta/2$. We consider a pentagonal defect with equilibrium tissue height of the hexagonal vertex model, i.e., $\tilde{h}=2^{1/3}\ 3^{-1/6}\ (2\Gamma)^{2/3}$.
@@
Fig. 7 shows the dimensionless core energy, Eq. (III.1), for different opening angles $\delta/2$ as a function of the apical edge length $a$ for (a) $\Gamma=\Gamma_{\mathrm{a}}=\Gamma_{\mathrm{b}}=0.65$ and (b) $1.0$, respectively. Edge length $a$ is bounded from above by the volume constraint in the unbulged vertex model case with $h_{\mathrm{S}}^{\mathrm{b}}=h_{\mathrm{S}}^{\mathrm{a}}=0$. For each angle the energy initially decreases if we decrease $a$ with concomitant outward bulging (top panel), as the area to volume ratio decreases. For even smaller $a$ the energy usually increases again, suggesting that a configuration with curved interfaces is always energetically more favorable.
@@
Increasing the angle leads to an increase in the core energy in the vertex model case (cf. dashed line at which the curves end in the top panel) and shifts the minimum of the energy toward smaller $a$ if we allow for bulging. If we compare the vertex model energy with the minimal energy in the bubbly vertex model for given opening angles (and thus surrounding tissue curvature), we find a smaller energy in the latter case, see lower panel.
@@ III.2 Deformation of the mean-field tissue
We use the description as a moderately bent plate as derived in Ref. [19]. The energy change due to the defect deformation can then be split into different contributions. First, the bending energy will change, both for the mean and Gaussian curvature contributions, due to changes in the curvature radii and opening angles in the inner cap segment. Second, the stretching energy will change as the azimuthal strain stemming from the pentagonal defect can be relaxed through changes in the curvature. For simplicity we will neglect radial stretching contributions at the interfaces by considering stress free boundary conditions at the contact points of defect cell, cap segment and outer cone segment and at the outer edge of the cone. This is justified numerically in Appendix F.
@@
The different energy contributions for bending and stretching for this mean-field tissue parameterization are derived in Appendix F. This parameterization with a mean-field model has been verified on icosahedral vertex model shells without interfacial curvatures, where the mentioned decomposition into a defect cell and a mean-field tissue is able to describe the tip much better than an approach purely based on continuum theory, cf. Appendix G.
@@
III.3 Extrusion of a bubbly defect cell in a mean-field vertex model tissue at icosahedral tips
We now consider a bubbly defect cell in a mean-field tissue of icosahedral symmetry, i.e., $h_{\mathrm{S}}^{\mathrm{b}}$ and $h_{\mathrm{S}}^{\mathrm{a}}$ may deviate from zero.